OpenStudy (swaqqout_kid1):

Ron plans on visiting an amusement park this weekend. Admission to the park costs $14.00, and each attraction costs $1.31 to use. If Ron has saved $64.00 for this weekend, what is the highest number of attractions he can visit at the park? 4 39 48 38 @sangya21 @Rachella @Angelina_Vella

OpenStudy (anonymous):

64.00 = 14.00 +1.31(x) x = no. of attraction 1.31(x) = 64.00 - 14.00 1.31(x) = 50.00 x = 50/1.31

OpenStudy (anonymous):

64-14=50 50/1.31= 38.17 so he can go to 31 rides